# Autonomous Horizon-based Asteroid Navigation With Observability-constrained Maneuvers

### v1

@@ Abstract
Asteroid exploration is a pertinent challenge due to the varying complexity of their dynamical environments, shape, and communication delays due to distance. Thus, autonomous navigation methods are continually being developed and improved in current research to enable their safe exploration. These methods often involve using Optical Navigation (OpNav) to determine the spacecraft’s location, which relies on the horizon’s visibility. It is critical to ensure the reliability of this measurement so that the spacecraft may maintain an accurate state estimate throughout its mission. This paper presents an observability-constrained algorithm that generates orbital maneuvers for spacecraft to follow trajectories that allow continuously usable optical measurements to maintain system observability for safe navigation. This algorithm improves upon existing asteroid navigation capabilities by allowing the safe and robust autonomous targeting of various trajectories. The algorithm is adaptable to different asteroid scenarios, and allows a spacecraft to orbit at a wide range of distances within the optical measurement range. The paper presents a comprehensive framework that simulates asteroid dynamics, synthetic image generation, edge detection, horizon-based OpNav, filtering, and observability-constrained control. [^1] A preliminary version of this work was previously presented during the 4th Space Imaging Workshop at Georgia Institute of Technology, Atlanta, GA.
@@
Asteroid Navigation, Horizon-based Optical Navigation, Extended Kalman Filtering, Lyapunov control, Spacecraft Autonomy, Observability-constrained control
@@ 1 Introduction
Asteroid exploration has recently gained significant interest, driven by scientific research, resource utilization, and planetary defense purposes. Successful missions, such as Hayabusa2 to 162173 Ryugu and OSIRIS-REx to 101955 Bennu, have provided valuable insights into the composition and structure of these small bodies [1, 2]. However, both missions required extensive observation periods to safely navigate their respective targets’ uncertain and weak gravitational environments. For example, OSIRIS-REx spent nearly a year mapping Bennu before attempting proximity operations [3], while Hayabusa2 similarly mapped Ryugu for months [4]. In addition to these missions, the DART and Lucy missions also emphasize the need for detailed pre-mission observations to ensure safe proximity operations [5, 6]. These missions highlight the need for improvements in autonomous Optical Navigation (OpNav) to reduce the reliance on long observation periods or human-in-the-loop processes to improve mission efficiency and reliability.
OpNav has become a preferred method for space missions due to its reliability, accuracy, and accessibility, particularly when navigating larger spherical bodies like the Moon or Mars [7]. OpNav relies on visual images of celestial bodies, allowing spacecraft to determine their position by analyzing surface features or horizons. Some types of OpNav include surface feature tracking, Line Of Sight (LOS)[8], Central and Apparent Diameter (CAD)[9], Lidar-based[10, 11], Limb-based [12] and Pole-from-Silhouette, Shape-from-Silhouette or Localization-from-Silhouette [13] methods.
In the case of Hayabusa2, both human-in-the-loop and autonomous OpNav are used extensively throughout the mission, from approach to landing and sample collection phases at asteroid 162173 Ryugu. The spacecraft employed a series of optical images captured by its onboard OpNav Camera (ONC) to estimate its position and velocity relative to the asteroid’s surface. This information is crucial for accurately guiding the spacecraft during its low-altitude descents and surface interactions. In addition, OpNav is used during the mission’s descent to provide detailed images of Ryugu’s surface, enabling the mission team to select safe landing sites. The mission relied heavily on OpNav for proximity operations, including final descent and collection maneuvers, illustrating its importance in achieving mission success. It also used a marker system for accurate autonomous operation near the asteroid [14].
Similarly, the OSIRIS-REx mission to asteroid 101955 Bennu also utilized human-in-the-loop OpNav during its approach, orbital phases, and sample collection. The spacecraft’s onboard navigation system analyzed optical images captured by its navigation camera to determine its position relative to Bennu, particularly during its low-altitude maneuvers. This data allowed OSIRIS-REx to navigate Bennu’s complex gravitational field, helping it to select sample collection sites and safely perform touch-and-go operations. The OpNav system provided real-time data during the mission’s critical phases, including orbital insertion, detailed surface mapping, and final sample collection. Throughout the mission, OpNav proved invaluable in guiding the spacecraft during autonomous operations when Earth-based communication delays are too long for manual intervention [15].
Existing OpNav methods have limitations that make them insufficient for autonomous spacecraft navigation. Natural feature tracking OpNav [15] is computationally expensive and may not be effective at large distances or near the dark side while using a physical marker method [14] requires close proximity to the asteroid for marker deployment . LOS [8] is advantageous due to its ease of use and applicability at large distances but at the cost of accuracy without range information. The newer silhouette methods [13] offer promising accuracy and determine asteroid poles and shapes in addition to localization but may be limited by computation time, except LimbNav [12] which is sufficiently computationally efficient for usage onboard spacecraft.
Another type of OpNav involves the horizon-based method. Unlike the other aforementioned methods, horizon-based OpNav measures the apparent horizon of the celestial body to determine the spacecraft’s position and orientation, making it effective at a wide range of distances and lighting conditions [16]. Furthermore, it is computationally efficient and feasible for real-time autonomous applications. Horizon-based OpNav has yet to be fully utilized in asteroid missions, but its proven accuracy and reliability in spherical-body missions suggest it could be adapted to advance asteroid navigation capabilities. By using the asteroid’s horizon rather than relying on surface features, horizon-based OpNav could ensure continuous and accurate navigation. It expands the range and angles at which measurements are sufficiently accurate to be usable and is effective from perspectives with low surface visibility, such as near the dark side. If successfully adapted for asteroid missions, it could help improve spacecraft autonomy, early mission approach time efficiencies, and reduce the need for manual interventions during critical mission phases. However, it is inaccurate when observing directly aligned with the lit or dark sides, as the horizon is detected erroneously or invisible.
Therefore, there is a need to develop a computationally efficient, autonomous, robust, and accurate OpNav and control algorithm. Ideally, this algorithm should function at a wide range of distances and angles while accounting for regions with poor observability to ensure continuous and safe operation. Various studies discuss robust orbit control around asteroids [17, 18, 19, 20, 21, 22, 23, 24] but do not explicitly consider the navigation performance in their formulations. This research proposes a novel solution: developing a path-constrained Lyapunov controller that uses constraints to avoid poor optical observability during asteroid missions using the horizon-based method. Lyapunov control has been explored for application in the asteroid mission context, however the focus has been on orbit-attitude control and hovering operations [25, 26, 27]. This newly proposed controller is dedicated to improving the OpNav measurement method and ensuring that the spacecraft avoids regions with poor visibility by adjusting its trajectory in real-time. This approach allows for continuous and reliable optical measurements, even near the asteroid’s dark side, thereby enhancing the overall safety and robustness of autonomous asteroid navigation.
@@
There are existing techniques that enhance OpNav performance using different approaches than the one discussed in this paper. [28] introduces an approach to optimize the spacecraft’s orbit during the approach phase by enhancing OpNav observability using the Fisher Information Matrix (FIM). While this method significantly improves measurement reliability, it is still challenged by uncertainties due to the asteroid’s irregular surface and dynamic environment, which could introduce errors. Additionally, the optimization process is too time-consuming and computationally costly, thus posing a challenge for its onboard application. In another work [29], the authors develop data-driven image processing techniques for OpNav to enhance navigation accuracy around binary asteroids. However, the computational demand for processing large datasets in real-time presents a challenge, potentially limiting onboard resources. [30] presents an observability-based navigation strategy that combines optical and radiometric measurements to improve navigation during the asteroid approach phase. The spacecraft configuration is optimized to enhance measurement accuracy and state estimation. However, this is only applicable to multi-spacecraft formations.
@@
The core technical approach of this paper involves deriving a Lyapunov controller that utilizes artificial potential functions to maintain a path with optimal observability. Additionally, we develop a synthetic asteroid image generation tool to simulate varying lighting conditions and test the performance of horizon-based OpNav in different scenarios. These simulations allow us to identify the limitations of existing OpNav methods, which then serve as inputs to our controller design. The proposed controller is tested in simplified mission scenarios, including orbit maintenance and approach for asteroid capture, with state estimation handled using an Extended Kalman Filter (EKF). The controller’s performance is evaluated using Monte Carlo simulations to ensure robustness across various mission profiles. A schematic displaying the flow of the algorithm is shown in Fig. 1
@@
The remainder of this paper is structured as follows. First, we discuss the dynamics model, horizon-based OpNav, and EKF, which form the foundation of our navigation system. Next, we describe the asteroid image generation technique and the edge detection methods used to test OpNav and analyze its limitations. Following this, we present the novel contribution of the observability-constrained Lyapunov controller. Finally, we demonstrate the controller’s effectiveness by analyzing OpNav and EKF performance, validated using Monte Carlo simulations under simplified mission conditions.
@@ 2.1 Dynamics
The motion of a spacecraft near primitive celestial bodies, such as asteroids, is influenced by several forces, including the body’s gravitational pull and the Sun’s Solar Radiation Pressure (SRP). The equations governing this motion are encapsulated in the Augmented Normalized Hill Three-Body Problem (ANH3BP), which provides a normalized framework accounting for these forces in a Cartesian rotating reference frame.
The Hill frame is defined with the $x$-axis pointing from the Sun to the asteroid, the $z$-axis pointing toward the angular velocity of the asteroid, and the $y$-axis completing the frame.
@@
where $G_{1}$ is the solar flux constant and $m/A$ is the spacecraft mass-to-area ratio. It is important to note that this model involves approximation due to the assumptions that the asteroid is in a circular orbit and SRP acceleration is constant.
@@
An alternate representation of these dynamics can be formulated using Milankovitch elements. First, let ${}^{H}\bm{x}_{\mathrm{Milan}}$ be the state of our system in Milankovitch elements and express the equation of motion as follows:
@@
where the inertial rate of change of the angular momentum vector and eccentricity vector are $\dot{\bm{h}}$ and $\dot{\bm{e}}$ respectively. $\bm{a}_{d}$ includes the perturbing accelerations of SRP, solar gravity and the control input. The rotation terms resolve to $-\tilde{\bm{\Omega}}\bm{h}=[-\Omega h_{1},\Omega h_{2},0]^{\top}$ and $-\tilde{\bm{\Omega}}\bm{e}=[-\Omega e_{1},\Omega e_{2},0]^{\top}$. $\bm{h}=\tilde{\bm{r}}\bm{v}$ is the angular momentum vector, $\bm{e}$ is the eccentricity vector and the asteroid’s angular momentum vector is $\bm{\Omega}=[0,0,\Omega]$. The tilde operator notation is used on 3D vectors to create the cross product matrix in the following form:
@@
${}^{H}\bm{a}_{\mathrm{Solar}}$ $=-\mu\left(\frac{{}^{H}\bm{r}}{\left\|{}^{H}\bm{r}\right\|_{2}^{3}}+\frac{{}^{H}\bm{r}_{\mathrm{asteroid}}}{\left\|{}^{H}\bm{r}_{\mathrm{asteroid}}\right\|_{2}^{3}}\right)$
where ${}^{H}\hat{\bm{z}}$ is the unit vector in the z-axis direction in the Hill frame, $h$ is the angular momentum magnitude, $h_{1}$ and $h_{2}$ are the angular momentum components in the x and y directions, $e_{1}$ and $e_{2}$ are the eccentricity components in the x and y directions, and ${}^{H}\bm{r}_{\mathrm{asteroid}}=\left[R,0,0\right]^{\top}$ in the Hill frame.

### v2

# Autonomous Observability-constrained Lyapunov Control for Horizon-based Small Body Navigation

@@ Abstract
Small body exploration is a pertinent challenge due to low gravity environments and strong sensitivity to perturbations like Solar Radiation Pressure (SRP). Thus, autonomous methods are being developed to enable safe navigation and control around small bodies. These methods often involve using Optical Navigation (OpNav) to determine the spacecraft’s location. Ensuring OpNav reliability would allow the spacecraft to maintain an accurate state estimate throughout its mission. This research presents an observability-constrained Lyapunov controller that steers a spacecraft to a desired target orbit while guaranteeing continuous OpNav observability. We design observability path constraints to avoid regions where horizon-based OpNav methods exhibit poor performance, ensuring control input that maintains good observability. This controller is implemented with a framework that simulates small body dynamics, synthetic image generation, edge detection, horizon-based OpNav, and filtering. We evaluate the approach in two representative scenarios—orbit maintenance and approach with circularization—around spherical and ellipsoidal target bodies. In Monte Carlo simulations, the proposed approach improves the rate of attaining target orbits without observability violations by up to 94% compared to an unconstrained Lyapunov baseline, demonstrating improved robustness over conventional methods.  [^1] A preliminary version of this work was previously presented during the 4th Space Imaging Workshop at Georgia Institute of Technology, Atlanta, GA.
@@
Small Body Navigation, Horizon-based Optical Navigation, Extended Kalman Filtering, Spacecraft Autonomy, Observability-constrained Lyapunov control
@@ 1 Introduction
Small body exploration has gained significant interest, driven by scientific research, resource utilization, and planetary defense purposes. Successful missions, such as Hayabusa2 to 162173 Ryugu and OSIRIS-REx to 101955 Bennu, have provided valuable insights into the composition and structure of these small bodies [1, 2]. However, both missions used human-in-the-loop OpNav throughout the mission and required extensive observation periods to safely navigate their respective targets’ uncertain and weak gravitational environments, which pose a significant navigational challenge due to trajectories being highly sensitive to perturbations like Solar Radiation Pressure (SRP) that may cause spacecraft deviation into poor observability regions or unsafe paths[3, 4]. For example, OSIRIS-REx spent nearly a year mapping Bennu before attempting proximity operations [5], while Hayabusa2 similarly mapped Ryugu for 3 months [6]. In addition to these missions, the DART and Lucy missions also emphasize the need for detailed pre-mission observations to ensure safe proximity operations [7, 8]. These missions highlight the need for improvements in autonomous Optical Navigation (OpNav) to reduce the reliance on long observation periods or human-in-the-loop processes to improve mission efficiency and reliability.
OpNav has become a preferred method for space missions due to its reliability, accuracy, and accessibility, particularly when navigating larger spherical bodies like the Moon or Mars [9]. OpNav relies on visual images of celestial bodies, allowing spacecraft to determine their position by analyzing surface features or horizons. Some types of OpNav include surface feature tracking, Line Of Sight (LOS) [10], Central and Apparent Diameter (CAD) [11], Lidar-based [12, 13], Limb-based [14] and Pole-from-Silhouette, Shape-from-Silhouette or Localization-from-Silhouette [15] methods. The type of OpNav used in this research is the Christian-Robinson Algorithm (CRA). The CRA is a type of horizon-based OpNav method that calculates the position of an observer with respect to the center of mass of an ellipsoid target body directly from the edge-detected apparent horizon points in an image rather than using curve fitting, and it has been used extensively in spherical-body missions and shown to be computationally efficient, accurate and feasible for real-time autonomous applications [16]. However, like other horizon-based methods, CRA’s performance strongly depends on the viewing geometry and lighting conditions. When the spacecraft is positioned near the dark side or in configurations with limited visible limb curvature, the quality of the measurements can degrade sharply due to poor observability, reducing the ability of the navigation filter to maintain accurate state estimates. Ensuring persistent observability throughout a mission is therefore a critical challenge for autonomous navigation around small bodies.
@@
There are existing techniques that maintain OpNav observability, however, they are computationally expensive and infeasible for autonomous application onboard a single spacecraft. [17] introduces an approach to optimize the spacecraft’s orbit during the approach phase by enhancing OpNav observability using the Fisher Information Matrix (FIM). While this method significantly improves measurement reliability, it is still challenged by uncertainties due to the asteroid’s irregular surface and dynamic environment, which could introduce errors. Additionally, the optimization process is too time-consuming and computationally costly, thus posing a challenge for its onboard application. In another work [18], the authors develop data-driven image processing techniques for OpNav to enhance navigation accuracy around binary asteroids. However, the computational demand for processing large datasets in real-time presents a challenge, potentially limiting onboard resources. [19] presents an observability-based navigation strategy that combines optical and radiometric measurements to improve navigation during the asteroid approach phase. The spacecraft configuration is optimized to enhance measurement accuracy and state estimation. However, this is only applicable to multi-spacecraft formations.
Small body proximity operations demand controllers that can satisfy safety and mission constraints such as keep-out zones, actuator limits, field-of-view/lighting requirements for navigation sensors while remaining computationally feasible for onboard usage. Optimization-based methods have been developed extensively towards this demand. Convex programming approaches can include thrust bounds, approach corridors, and state-triggered constraints with formal feasibility guarantees. Some examples include lossless convexification and successive convexification for powered descent and proximity guidance, which deliver real-time, fuel-efficient solutions under tight constraints [20, 21, 22]. Another type is Model Predictive Control (MPC), which has been demonstrated for autonomous rendezvous and docking with explicit state/control constraints and keep-out regions, achieving robust performance against target tumbling and environmental uncertainties [23, 24, 25]. Complementary safety-critical control with control barrier functions (CBFs) provides a Lyapunov-like invariance mechanism enforced through online quadratic programs, now a matured tool to guarantee set invariance (e.g., collision avoidance, line-of-sight cones) while permitting separate performance objectives [26, 27]. On the estimation side, observability has also been treated as a constraint or objective since angles-only navigation is known to lose range observability along certain trajectories thereby motivating information-aware trajectory design and “observability maneuvers” that enforce good view-geometry and lighting constraints [28, 29, 30, 31]. This paper builds upon similar principles by treating navigation observability and sensing geometry as operational constraints to be considered by the controller through path constraints.
Lyapunov methods offer nonlinear, feedback designs with provable stability and light onboard computational cost, which is ideal for this application. These have seen sustained use across space applications in the form of hovering and precision descent near small bodies via sliding-mode controllers with Lyapunov proofs [32, 33], formation flying and relative motion regulation using orbital-element feedback and artificial potentials [34, 35], and attitude/pose regulation under uncertainties [36]. Classical Lyapunov designs target asymptotic convergence without explicitly encoding path constraints. However, constraint handling can be integrated by augmenting the Lyapunov construction with artificial potentials, state penalties, or by coupling with barrier-type invariance conditions, which is an approach that bridges to the constrained Lyapunov and CBF literature [26, 27]. Recent advances even address constrained Lyapunov stabilization for orbital transfer problems using Gauss variational equations, illustrating how Lyapunov functions and constraint satisfaction can coexist within a principled design [37]. Taken together, this literature shows Lyapunov control is both established for nonlinear spacecraft dynamics and extensible to include constraint-aware control.
Due to the limitations of current OpNav observability maintaining techniques and the controller demands of small body proximity operation applications, there is a need to develop a computationally efficient, robust, and accurate observability-maintaining control algorithm to facilitate safe, robust and autonomous small body navigation. Ideally, this algorithm should function at a wide range of distances and angles while accounting for regions with poor observability to ensure continuous and safe operation. Various studies discuss robust orbit control around asteroids. [38, 39] apply stochastic optimal control to optimize trajectories under gravitational and operational uncertainties, with more recent work employing chance-constrained convex programming for safe autonomy [40], chance-constrained sensing-optimal path planning for angles-only navigation [41], and sequential covariance steering for robust cislunar low-thrust trajectory optimization under uncertainty [42]. [43] uses state transition tensors for rapid autonomous orbit transfers after parameter estimation, [44, 45] develop uncertainty-aware rendezvous optimization, [46, 47] apply chance-constrained convex optimization for stochastic maneuver design, and [48] ensures stability through Lyapunov redesign under mass distribution uncertainties. However, they do not explicitly consider the navigation performance in their formulations.
This research proposes a novel solution: developing a Lyapunov controller that uses path constraints to avoid poor optical observability during asteroid missions using the CRA. Lyapunov control has been explored for application in the asteroid mission context, however the focus has been on orbit-attitude control and hovering operations [49, 50, 51]. This newly proposed controller is dedicated to maintaining the usability of OpNav by ensuring that the spacecraft avoids regions with poor observability through real-time trajectory adjustments. This approach allows for continuous and reliable optical measurements, even near the asteroid’s dark side, thereby enhancing the overall safety and robustness of autonomous asteroid navigation.
In the problem context of this paper, the controller must keep the spacecraft in observability-favorable configurations for OpNav, effectively a path constraint tied to lighting and viewing geometry, while remaining simple and robust enough for onboard execution. Optimization-centric schemes (convex/MPC) offer powerful guarantees but incur nontrivial online solve times, tuning/initialization overhead, and algorithmic complexity [21, 23, 24]. By contrast, a Lyapunov-based design yields a closed-form feedback law with stability guarantees and negligible per-step compute, and it can incorporate observability-related path constraints through smooth artificial potentials or barrier-like terms, which is conceptually aligned with safety-critical control [26, 27]. This makes Lyapunov control an appropriate choice to solve this problem as it preserves stability and respects observability-based constraints as documented in observability-aware guidance [28, 30, 31], and remains lightweight enough for continuous onboard operation during extended small-body proximity phases.
@@
The core technical approach of this paper involves deriving a Lyapunov controller that utilizes artificial potential functions to maintain a path with optimal observability. Additionally, we develop a synthetic image generation tool to simulate varying lighting conditions and test the performance of horizon-based OpNav in different scenarios. These simulations allow us to identify the limitations of existing OpNav methods, which then serve as inputs to our controller design. The proposed controller is tested in simplified mission scenarios, including orbit maintenance and approach for small body capture, with state estimation handled using an Extended Kalman Filter (EKF). The controller’s performance is evaluated using Monte Carlo simulations to ensure robustness across various mission profiles. A schematic displaying the flow of the algorithm is shown in Fig.˜1.
@@
The remainder of this paper is structured as follows. First, we discuss the dynamics model, horizon-based OpNav, and EKF, which form the foundation of our navigation system. Following this, we present the novel contribution of the observability-constrained Lyapunov controller. Finally, we demonstrate the controller’s effectiveness by analyzing OpNav and EKF performance, validated using Monte Carlo simulations under simplified mission conditions. The synthetic image generation pipeline and edge detection techniques used are described in appendix˜A.
@@ 2.1 Dynamics
The motion of a spacecraft near small bodies is influenced by several forces, including the body’s gravitational pull and the Sun’s Solar Radiation Pressure (SRP). The equations governing this motion are encapsulated in the Augmented Normalized Hill Three-Body Problem (ANH3BP), which provides a normalized framework accounting for these forces in a Cartesian rotating reference frame.
The Hill frame is defined with the $x$-axis pointing from the Sun to the small body, the $z$-axis pointing toward the angular velocity of the small body, and the $y$-axis completing the frame.
@@
where $G_{1}$ is the solar flux constant and $m/A$ is the spacecraft mass-to-area ratio. It is important to note that this model involves approximation due to the assumptions that the small body is in a circular orbit and SRP acceleration is constant.
@@
An alternate representation of these dynamics can be formulated using Milankovitch elements. Milankovitch elements are used for the controller because they have no singularity and can target a certain orbit state over a specific point. First, let $^{H}\bm{x}_{\mathrm{Milan}}$ be the state of our system in Milankovitch elements and express the equation of motion as follows:
@@
where the inertial rate of change of the angular momentum vector and eccentricity vector are $\dot{\bm{h}}$ and $\dot{\bm{e}}$ respectively. $\bm{a}_{d}$ includes the perturbing accelerations of SRP, solar gravity and the control input. The rotation terms resolve to $\tilde{\bm{\Omega}}\bm{h}=[-\Omega h_{1},\Omega h_{2},0]^{\top}$ and $\tilde{\bm{\Omega}}\bm{e}=[-\Omega e_{1},\Omega e_{2},0]^{\top}$. $\bm{h}=\tilde{\bm{r}}\bm{v}$ is the angular momentum vector, $\bm{e}$ is the eccentricity vector and the small body’s angular momentum vector is $\bm{\Omega}=[0,0,\Omega]$. The tilde operator notation is used on 3D vectors to create the cross product matrix in the following form:
@@
${}^{H}\bm{a}_{\mathrm{Solar}}$ $=-\mu\left(\frac{{}^{H}\bm{r}}{\left\|{}^{H}\bm{r}\right\|_{2}^{3}}+\frac{{}^{H}\bm{r}_{\mathrm{sb}}}{\left\|{}^{H}\bm{r}_{\mathrm{sb}}\right\|_{2}^{3}}\right)$
where $^{H}\hat{\bm{z}}$ is the unit vector in the z-axis direction in the Hill frame, $h$ is the angular momentum magnitude, $h_{1}$ and $h_{2}$ are the angular momentum components in the x and y directions, $e_{1}$ and $e_{2}$ are the eccentricity components in the x and y directions, and $^{H}\bm{r}_{\mathrm{sb}}=\left[R,0,0\right]^{\top}$ in the Hill frame.
@@ 2.1.1 Frozen Terminator Orbit
2.1.2 Point Mass Feasibility
Modeling the small body as a point mass is sufficient as the simulations conducted in this research are at relatively high distances from small masses where irregular gravity field effects become negligible. To justify this assumption, we compare the magnitudes of the two most major potential sources of perturbation, J2 and SRP, to determine which one is more significant. Assuming the considered smooth small body has the properties of Bennu and is a uniform–density triaxial ellipsoid of mass $M=7.329\times 10^{10}\,\mathrm{kg}$, radius $r=0.241\,\mathrm{km}$ and shape ratios $[a,b,c]=[1.1051,1.0769,1]$, the principal semi-axes are $a,b,c]_{\mathrm{dim}}=\frac{r}{abc^{1/3}}[a,b,c]=[0.25158,\;0.24520,\;0.22752]\,\mathrm{km} [$2, 55]. Thus, the J2 (second zonal harmonic) coefficient follows
$J_{2}\;=\;\frac{I_{z}-\tfrac{1}{2}(I_{x}+I_{y})}{M\,R_{\rm ref}^{2}}\;\approx\;3.41\times 10^{-2}.$ (14)
[…]
The maximum magnitude of the $J_{2}$ perturbation acceleration at a spacecraft distance $d_{\mathrm{sat}}$ occurs over the poles and for the closest distance encountered in this paper’s simulations, $1\,\mathrm{km}$, is
$a_{J_{2}}^{\max}(d_{\mathrm{sat}})\;=\;\frac{3\,J_{2}\,\mu\,R_{\rm ref}^{2}}{d_{\mathrm{sat}}^{4}}\quad\Longrightarrow\quad a_{J_{2}}^{\max}(1\,\mathrm{km})\approx 2.8906\times 10^{-11}\;\mathrm{km/s^{2}},$ (15)
where $\mu=4.8904\times 10^{-9}\,\mathrm{km^{3}/s^{2}}$ is the small body’s gravitational parameter.
By contrast, the SRP acceleration in this case is $\beta\approx 1.0243\times 10^{-10}\,\mathrm{km/s^{2}}$ according to Eq.˜2 and the dynamical parameters outlined in Table˜1. The gravitational acceleration at $r=1\,\mathrm{km}$ from the small body is $a_{g}=4.8904\times 10^{-9}\,\mathrm{km/s^{2}}$. Comparing, we find that at its maximum, the $J_{2}$ acceleration is $0.5911\%$ of the gravitational acceleration and $28.2202\%$ of the SRP. Note that for the majority of the trajectories simulated, the spacecraft is not perfectly aligned with the poles and is at further distances, thereby experiencing even lower $J_{2}$ effects while SRP remains constant. Therefore, we can conclude that the effect of J2 is secondary to SRP, which remains the most significant perturbation. Accordingly, higher order gravity-field irregularities are of even lower significance and would be adjusted for by a filtering algorithm considering an appropriate process noise level. However, for small bodies with larger mass, highly elliptic shape, or at farther distances from the Sun, the J2 perturbation may become more significant than SRP. Therefore, if required, extension to mascon or polyhedron models may be performed by adding these small perturbation terms to the ANH3BP dynamics similarly as the SRP by using the control‐influence matrix $B$, with no change to the core EKF or Lyapunov‐controller structure.